OpenStudy (anonymous):

Find dy/dx. 2x^2y^2=x^3+y^3

OpenStudy (anonymous):

\[(2x^2y^2)\prime=(2x^2)(y^2)\prime+(2x^2)\prime(y^2)\]

OpenStudy (anonymous):

@esam2 can u calculate this ?^

OpenStudy (anonymous):

i'm not that good at this...

OpenStudy (anonymous):

\[(y^2)'=?\]

OpenStudy (anonymous):

2y

OpenStudy (anonymous):

2yy'

OpenStudy (anonymous):

\[(2x^2)'=?\]

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

true

OpenStudy (anonymous):

now can u differentiate right side?

OpenStudy (anonymous):

this is Implicit Functions

OpenStudy (anonymous):

I'm very new to this concept, i need some help

OpenStudy (anonymous):

\[(x^3)'= ? \] \[(y^3)'=?\]

OpenStudy (anonymous):

3x 3y^2 y'?

OpenStudy (anonymous):

correct . but \[(x^3)'=3x^2\]

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

\[(2x^2y^2)'=(x^3)'+(y^3)'\] \[(2x^2)(2yy')+(4x)(y^2)=(3x^2)+(3y^2y')\]

OpenStudy (anonymous):

now find y'

OpenStudy (anonymous):

\[\frac{ 3x^2-4xy^2 }{ 4x^2y-3y^2 }\] ???

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thank you! :)